# CompMarkGS: Robust Watermarking for Compressed 3D Gaussian Splatting

### v5

@@ 5.3 Ablation Study
We embed the watermark into different anchor attributes (position, scaling, and offsets) and compare the resulting rendering quality and bit accuracy. As discussed in the previous section (Sec.4.1), since the watermark embedding feature shares the same dimensions as the anchor feature, we ensure consistency across all attributes before embedding. As shown in Fig.4 and Tab.5, embedding the watermark into anchor attributes other than the anchor feature leads to degraded rendering quality compared to embedding it into the anchor feature.
In particular, the anchor position and the scaling lead to significant quality degradation because these attributes directly adjust the Gaussian pose and shape. These results demonstrate that our method seamlessly integrates the watermark into the scene’s features without significantly degrading the scene information.
Quantization distortion layer.
We evaluate the effectiveness of the quantization distortion layer (QDL) during the quantization process by comparing rendering quality and bit accuracy before and after compression, with and without QDL. We use HAC chen2024hac , a quantization-based 3DGS compression method. Tab.6 reports the results after compression for HAC, showing that performance without QDL is noticeably worse than with QDL. Without QDL during training, the anchor attributes carrying the watermark fail to adapt to the quantization process, leading to information loss and reduced bit accuracy.
Table 5: Ablation studies on watermark embedding into different anchor attributes. Results are reported for 48-bit watermarks and averaged over three datasets using HAC after compression.
Target Parameter Bit Acc(%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Position 67.91 12.05 0.625 0.350 Scaling 65.06 19.35 0.615 0.323 Offsets 94.88 24.61 0.792 0.226 Anchor Feature (Ours) 95.92 27.65 0.852 0.177
Table 6: Ablation studies on quantization distortion layer (QDL) and FAG. Based on 48-bit, results are averaged over three datasets with HAC after compression.
Methods  Bit Acc(%)↑  PSNR↑  SSIM↑  LPIPS↓ w/o QDL +  w/  FAG 90.75 26.75 0.844 0.182  w/  QDL + w/o FAG 93.95 27.54 0.849 0.178  w/  QDL +  w/  FAG 95.92 27.65 0.852 0.177
In contrast, applying QDL allows the anchor attributes to adapt to the quantization process during training, thereby minimizing information loss. These results demonstrate that the watermark embedded using QDL remains robust under quantization.
Frequency-aware anchor growing and HSV loss.
To evaluate our proposed frequency-aware anchor growing (FAG) and HSV loss $\mathcal{L}_{hsv}$ for mitigating distortion during watermark embedding, we compare performance with and without each component. As shown in Tab.6 and Fig.5, bit accuracy improves when FAG is enabled. This improvement arises because FAG selectively grows anchors in regions lacking detailed Gaussians, increasing the number of Gaussians available for watermark insertion. Additionally, rendering quality improves further than without HSV loss. Fig.5 also shows that HSV loss effectively removes color artifacts of rendered images. These results underscore the importance of both components in reducing watermark-induced distortions.
@@ 6 Conclusion
We propose CompMarkGS, a compression-robust watermarking method. By integrating a quantization distortion layer, we train the watermark to withstand quantization-based compression, and we introduce a learnable watermark embedding feature for anchor-based insertion, achieving both high security and high fidelity. Additionally, our frequency-aware anchor growing and HSV loss preserve high-quality rendering performance. To the best of our knowledge, CompMarkGS is the first compression-robust watermarking framework for 3DGS, making it well suited for protecting 3D assets on resource-constrained devices.
Limitations and future works.
CompMarkGS requires a pre-trained decoder. However, since the decoder needs training only once per bit-length, subsequent usage is efficient without additional fine-tuning. We will extend our method beyond 3DGS compression and develop a lightweight watermarking approach for 3DGS.
@@ Appendix A Broader impacts
Appendix B More implementation details
B.1 Details of training
@@
Our proposed CompMarkGS trains the anchor-based 3DGS from scratch with the watermark, in contrast to existing methods [13, 14, 15] that rely on fine-tuning pre-trained 3DGS models. We train CompMarkGS for $30,000$ iterations. Anchor growing is performed between $1,600$ and $15,000$ iterations, first using conventional anchor growing and then switching to the proposed frequency-aware anchor growing strategy for fine-grained refinement. The loss function for watermark training consists of the original anchor-based 3DGS [25] loss $\mathcal{L}_{\text{scaffold}}$ and three additional components: HSV loss $\mathcal{L}_{\text{hsv}}$, frequency loss $\mathcal{L}_{\text{freq}}$, and message loss $\mathcal{L}_{\text{msg}}$. Before watermark training begins, we pre-train a HiDDeN [56] message decoder on the MS-COCO dataset [24] for 32, 48, and 64-bit messages following the strategy from 3D-GSW [15] and Stable Signature [10]. During the subsequent watermark training, the pre-trained decoder is kept fixed. Additionally, the compression ratio of the compression method used in all experiments is set to $0.004$.
B.2 Details of high-pass filter
@@
where $d(p)$ denotes the Euclidean distance between a pixel $p$ and the center of the image. $\tau$ defines the cutoff threshold of the mask and $\beta$ determines the degree of attenuation. The resulting mask $M_{h}$ is multiplied element-wise with the Fourier-transformed image to suppress low-frequency components and preserve only the high-frequency details. Afterward, we apply the Inverse Discrete Fourier Transform (IDFT) to reconstruct the filtered images in the spatial domain, resulting in the high-frequency rendered images $I^{\prime}_{hf}$ and the high-frequency ground truth images $I_{hf}$. Further details of the frequency-aware anchor growing process can be found in the main paper. (See main paper Sec.4.3)
B.3 Details of frequency-aware anchor growing
We propose a frequency-aware anchor growing strategy that efficiently identifies and densifies Gaussians in high-frequency regions to mitigate visual quality degradation caused by watermark embedding. Specifically, we first compute a pixel-wise SSIM-based error map $P_{error}$ (See main paper Sec.4.3) between the high-frequency rendered image $I^{\prime}_{hf}$ and the ground truth image $I_{hf}$, both extracted in the frequency domain.(B.2) This error map $P_{error}$ highlights pixel regions where high-frequency information is not faithfully reconstructed. To identify pixel regions that contain visually important fine details, we compute the median value $\tilde{P}_{error}$ of the SSIM-based pixel-wise error map $P_{error}$. Using the median value $\tilde{P}_{error}$ ensures a more meaningful identification of high-frequency regions by mitigating the influence of outlier pixels, which may otherwise dominate when relying on the maximum value of the SSIM-based pixel-wise error map $P_{error}$. We then generate a binary mask $I_{mask}$ by filtering pixels whose error values fall within the interval $[\tilde{P}_{error}-\epsilon,\tilde{P}_{error}+\epsilon]$:
@@ B.4 Details of HSV loss
Although the RGB color space is widely used for image representation, prior work on image sharpening [17] indicates that the RGB color space does not fully account for the perceptual characteristics of the human visual system. In particular, the human visual system is more sensitive to luminance variations than to chromatic changes [11]. Unlike the RGB color space, the Hue, Saturation, and Value (HSV) space separates chromatic information (Hue and Saturation) from luminance (Value), enabling more precise detection and suppression of color artifacts. Based on this property, we introduce an HSV loss that enhances rendering quality while preserving high bit accuracy.
@@
$H_{R}\in\left[0,\dfrac{\pi}{3}\right)\cup\left[\dfrac{5\pi}{6},2\pi\right),\;\;\text{if}\;\;S(p)\geq\tau_{R}^{s}\;\;\text{and}\;\;V(p)\geq\tau_{R}^{v},$ (15)
@@
Appendix C Computing resources
We evaluate the rendering performance of our method and baselines [13, 14, 15] by measuring FPS. As shown in Tab.7, our method exceeds the real-time rendering threshold of $30$ FPS across the Blender [29], LLFF [28], and Mip-NeRF 360 [3] datasets. These results demonstrate the practical applicability and rendering efficiency of our proposed CompMarkGS.
FPS↑ Methods Blender LLFF Mip-NeRF 360 WateRF [14] 199.30 19.31 41.98 GaussianMarker [13] 166.45 16.24 28.54 3D-GSW [15] 220.60 22.67 50.64 CompMarkGS 230.39 35.30 67.58
Appendix D Additional results
D.1 Embedding the watermark in anchor attributes
We evaluate the impact of embedding the watermark into different anchor attributes (position, scaling, offsets) by measuring the rendering quality and bit accuracy. As discussed in the main paper (See main paper Sec.4.1), the proposed watermark embedding feature is adjusted to match the dimensionality of each anchor attribute to ensure compatibility during insertion. As reported in Tab. 8, embedding the watermark into attributes other than the anchor feature results in a noticeable degradation in rendering quality. Unlike the anchor feature, which predicts Gaussian attributes via individual MLPs, position and scaling directly affect the position and shape of each Gaussian. Consequently, embedding into these attributes causes significant degradation in both rendering quality and bit accuracy. Offsets, which represent the relative position of Gaussians from anchor points within each voxel, yield better bit accuracy than position and scaling but still suffer from lower visual quality. These results confirm that the proposed method can reliably embed the watermark into the model without compromising the scene fidelity.
@@ D.1 Embedding the watermark in anchor attributes
Table 8: Ablation studies on watermark embedding into different anchor attributes. Results are reported for 48-bit messages and averaged over three datasets using ContextGS after compression.
@@
Target Parameter Bit Acc(%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Position 82.37 18.49 0.632 0.432 Scaling 64.68 13.40 0.466 0.474 Offsets 93.98 25.32 0.755 0.322 Anchor Feature (Ours) 94.03 27.55 0.844 0.173
D.2 Effectiveness of our contributions
To evaluate the individual contribution of each component, we conduct ablation studies by selectively removing the frequency-aware anchor growing (FAG), quantization distortion layer (QDL), and HSV loss $\mathcal{L}_{hsv}$ from our proposed method. We perform under HAC [7] and ContextGS [39] compression settings, and we compare the resulting bit accuracy and rendering quality. As shown in Tab. 9, removing QDL causes a notable decrease in bit accuracy, demonstrating its effectiveness in preserving watermark information embedded in anchor attributes under quantization-based compression. In contrast, removing FAG maintains high bit accuracy but leads to degraded rendering quality. This highlights the role of FAG in enhancing visual fidelity by promoting anchor densification in high-frequency regions. Removing the HSV loss also results in a slight degradation in rendering quality, particularly in LPIPS. These results indicate its effectiveness in suppressing color artifacts that are not adequately addressed by RGB-based losses alone. Notably, we observe that removing all components results in the most significant performance degradation. These findings collectively demonstrate that FAG, QDL, and HSV loss each contribute critically to balancing bit accuracy and rendering quality under compression. The integration of all three components yields the best overall performance in compressed settings.
HAC+CompMarkGS ContextGS+CompMarkGS FAG QDL $\mathcal{L}_{hsv}$ Bit Acc(%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Bit Acc(%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ 90.67 26.44 0.827 0.196 87.63 25.56 0.812 0.223 $\checkmark$ $\checkmark$ 93.95 27.54 0.849 0.178 90.35 27.15 0.841 0.198 $\checkmark$ $\checkmark$ 90.75 26.75 0.844 0.182 88.09 25.04 0.801 0.229 $\checkmark$ $\checkmark$ 92.57 27.49 0.852 0.179 90.23 27.19 0.841 0.200 $\checkmark$ $\checkmark$ $\checkmark$ 95.92 27.65 0.852 0.177 94.03 27.55 0.844 0.173
D.3 Effectiveness of HSV loss
When only RGB-based loss functions are used for watermark embedding, the optimization of 3DGS fails to consider the characteristics of the human visual system. As a result, 3DGS with a watermark renders color artifacts in the rendered images, particularly in regions with strong chromatic and luminance variations. These artifacts are observed in Fig.6 and Fig.7, despite achieving high bit accuracy.

### v6

@@ 5.3 Ablation Study
We embed the watermark into different anchor attributes (position, scaling, and offsets) and compare the rendering quality and bit accuracy. As discussed in the previous section (Sec.4.1), since the watermark embedding feature shares the same dimensions as the anchor feature, when we embed the watermark into a different attribute, we set the watermark feature to match the dimensions of that specific attribute for the experiment. As shown in Fig.5 and Tab.4, embedding the watermark into anchor attributes such as position and scaling leads to significant quality degradation, as these attributes directly adjust the Gaussian pose and shape. In contrast, our method embeds the watermark into the anchor feature, seamlessly concealing it without significantly degrading scene information.
Effectiveness of our contributions.
We conduct an ablation study to analyze the individual contribution of our key components: frequency-aware anchor growing (FAG), quantization distortion layer (QDL), and the HSV loss $\mathcal{L}_{hsv}$. The experiments are performed under HAC and ContextGS compression settings, with the results reported in Tab.5. Removing QDL causes the most significant drop in bit accuracy, demonstrating its crucial role in preventing watermark loss during quantization. Removing FAG degrades both rendering quality and bit accuracy. This indicates that by growing anchors exclusively in high-frequency regions, FAG mitigates conflicts between low-frequency watermarks and quality enhancement algorithms, thereby securing both rendering quality and watermark robustness (See Fig.6). Finally, removing the HSV loss worsens the LPIPS score and introduces subtle color artifacts. This suggests that the HSV loss effectively mitigates color artifacts that are difficult to handle with an RGB-space loss function alone (See Fig.6). Effective watermarking must strike a balance among three competing goals: robustness, invisibility, and capacity. Our experiments show that FAG, QDL, and the HSV loss not only serve unique roles but also work complementarily to significantly improve this balance between bit accuracy and rendering quality in compressed environments.
@@ 6 Conclusion
We propose CompMarkGS, a compression‐robust watermarking method. By integrating a quantization distortion layer, we train the watermark to withstand quantization‐based compression, and we introduce a learnable watermark embedding feature for anchor‐based insertion, achieving both high security and high fidelity. Additionally, our frequency‐aware anchor growing and HSV loss preserve high‐quality rendering performance. To the best of our knowledge, CompMarkGS is the first compression‐robust watermarking framework for 3DGS, making it well-suited for protecting 3D assets on resource‐constrained devices.
Ethics Statement
Provenance.
We only used publicly released datasets from academic publications (Blender (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022)). Additionally, the Blender and LLFF datasets are distributed under the CC BY 3.0 Unported license, while the Mip-NeRF 360 dataset is available under the Apache License 2.0.
Anonymous.
The Blender (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022) datasets represent synthetic and real-world data, respectively. The Blender dataset consists of 3D object scenes, while LLFF and Mip-NeRF 360 comprise indoor and outdoor real-world scenes. None of these datasets contains any human subjects. The details of each dataset are as shown in Tab.6.
Table 6: Dataset composition used for evaluation. Blender is synthetic, whereas LLFF and Mip-NeRF 360 are real-world scenes, with no human subjects included.
Dataset Scene1 Scene2 Scene3 Scene4 Scene5 Scene6 Scene7 Scene8 Scene9 Blender chair drums ficus hotdog lego materials mic ship LLFF fern flower fortress horns leaves orchids room trex Mip-NeRF 360 bicycle bonsai counter flowers garden kitchen room stump treehill
The use of large language models
We used an LLM in a limited manner for grammar and style polishing only. It did not contribute to research ideation, experimental design, analysis. The authors take full responsibility for all content.
@@ Appendix A Broader impacts
Appendix B Limitations and future works.
B.1 Limitations
One limitation of our proposed method, CompMarkGS, is its reliance on a pre-trained decoder for the watermark extraction process. However, this dependency constitutes a manageable, one-time computational cost. The decoder is trained only once for a specific message length (e.g., 32, 48, or 64 bits) and can then be universally applied to any 3D Gaussian Splatting model without the need for model-specific fine-tuning. This design choice significantly enhances the method’s practicality and scalability, as the initial training effort is amortized over countless applications.
B.2 Future works
Looking ahead, we have identified several promising directions for future work. First, we plan to extend our method to operate independently of 3D Gaussian Splatting compression, which would broaden its applicability to a wider range of scenarios. Concurrently, we aim to develop a more lightweight watermarking architecture. This initiative will focus on reducing the computational overhead of both the embedding and decoding stages, making our solution even more suitable for real-time or resource-constrained environments.
Appendix C More implementation details
C.1 Target Compression Schemes for Robustness Evaluation
To test the robustness of our watermark, we use two prominent anchor-based 3DGS compression methods HAC (Chen et al., 2024a) and ContextGS (Wang et al., 2024a). Both techniques follow a similar pipeline, extracting features for each anchor, using MLPs to predict a context-based distribution, and then applying differentiable quantization and entropy coding. This approach significantly compresses the model size while minimizing any loss in rendering quality.
However, the information loss that occurs during quantization can damage or destroy a watermark. Therefore, our core objective is to achieve robustness specifically against this kind of quantization-based compression.
C.2 Overview of the CompMarkGS pipeline
In this section, we detail the core components of the CompMarkGS pipeline, covering the overall watermark training procedure.
Watermarking embedding stage.
The watermarking process begins by combining the original anchor features with a learnable watermark embedding feature to create watermarked anchor features. These features are then passed through a quantization distortion layer (QDL), which simulates distortions from quantization by injecting noise, resulting in the quantized watermarked anchor features. These final features are fed into the MLPs to predict Gaussian parameters and render the scene. To recover the watermark message from the rendered image, we apply a Discrete Wavelet Transform (DWT) to extract the low-frequency band and input it into a pre-trained decoder. The entire training process is optimized to minimize the message reconstruction loss $\mathcal{L}_{msg}$, with an additional HSV-based perceptual loss $\mathcal{L}_{hsv}$ used to prevent visual quality degradation.
Frequency-aware anchor growing stage.
To further enhance rendering quality, we introduce the frequency-aware anchor growing (FAG) stage, which leverages frequency information to grow new anchors. We apply a Discrete Fourier Transform (DFT) and a high-pass filter to both the rendered and ground truth images to isolate their high-frequency components. A pixel-wise SSIM error map is then calculated between these two high-frequency representations. Based on this error map, we identify Gaussians in regions where high-frequency information is poorly reconstructed and use backpropagation to selectively grow new anchors in those locations.
In summary, CompMarkGS learns watermark robustness against compression-induced information loss via the QDL, while simultaneously preserving high-frequency details that could be degraded by watermarking via FAG. As a result, our proposed method achieves high watermark recovery rates and excellent visual fidelity in scenarios both with and without model compression.
C.3 Details of training
@@
Our proposed CompMarkGS trains the anchor-based 3DGS from scratch with the watermark, in contrast to existing methods Huang et al. (2025); Jang et al. (2024; 2025) that rely on fine-tuning pre-trained 3DGS models. We train CompMarkGS for $30,000$ iterations. Anchor growing is performed between $1,600$ and $15,000$ iterations, first using conventional anchor growing and then switching to the proposed frequency-aware anchor growing strategy for fine-grained refinement. The loss function for watermark training consists of the original anchor-based 3DGS Lu et al. (2024) loss $\mathcal{L}_{\text{scaffold}}$ and three additional components: HSV loss $\mathcal{L}_{\text{hsv}}$, frequency loss $\mathcal{L}_{\text{freq}}$, and message loss $\mathcal{L}_{\text{msg}}$. Before watermark training begins, we pre-train a HiDDeN Zhu et al. (2018) message decoder on the MS-COCO dataset Lin et al. (2014) for 32, 48, and 64-bit messages following the strategy from 3D-GSW Jang et al. (2025) and Stable Signature Fernandez et al. (2023). During the subsequent watermark training, the pre-trained decoder is kept fixed. Additionally, the compression ratio of the compression method used in all experiments is set to $0.004$.
C.4 Details of high-pass filter
@@
where $d(p)$ denotes the Euclidean distance between a pixel $p$ and the center of the image. $\tau$ defines the cutoff threshold of the mask and $\beta$ determines the degree of attenuation. The resulting mask $M_{h}$ is multiplied element-wise with the Fourier-transformed image to suppress low-frequency components and preserve only the high-frequency details. Afterward, we apply the Inverse Discrete Fourier Transform (IDFT) to reconstruct the filtered images in the spatial domain, resulting in the high-frequency rendered images $I^{\prime}_{hf}$ and the high-frequency ground truth images $I_{hf}$. Further details of the frequency-aware anchor growing process can be found in the main paper (See main paper Sec.4.3).
C.5 Details of frequency-aware anchor growing
We propose a frequency-aware anchor growing strategy that efficiently identifies and densifies Gaussians in high-frequency regions to mitigate visual quality degradation caused by watermark embedding. Specifically, we first compute a pixel-wise SSIM-based error map $P_{error}$ (See main paper Sec.4.3) between the high-frequency rendered image $I^{\prime}_{hf}$ and the ground truth image $I_{hf}$, both extracted in the frequency domain (See Sec.C.4). This error map $P_{error}$ highlights pixel regions where high-frequency information is not faithfully reconstructed. To identify pixel regions that contain visually important fine details, we compute the median value $\tilde{P}_{error}$ of the SSIM-based pixel-wise error map $P_{error}$. Using the median value $\tilde{P}_{error}$ ensures a more meaningful identification of high-frequency regions by mitigating the influence of outlier pixels, which may otherwise dominate when relying on the maximum value of the SSIM-based pixel-wise error map $P_{error}$. We then generate a binary mask $I_{mask}$ by filtering pixels whose error values fall within the interval $\tilde{P}_{error}-\epsilon,\tilde{P}_{error}+\epsilon]$:
@@ C.6 Details of HSV loss
Although the RGB color space is widely used for image representation, prior work on image sharpening Kau & Lee (2013) indicates that the RGB color space does not fully account for the perceptual characteristics of the human visual system. In particular, the human visual system is more sensitive to luminance variations than to chromatic changes Gonzalez (2009). Unlike the RGB color space, the Hue, Saturation, and Value (HSV) space separates chromatic information (Hue and Saturation) from luminance (Value), enabling more precise detection and suppression of color artifacts. Based on this property, we introduce an HSV loss that enhances rendering quality while preserving high bit accuracy.
@@
$H_{R}\in\left[0,\dfrac{\pi}{3}\right)\cup\left[\dfrac{5\pi}{3},2\pi\right),\;\;\text{if}\;\;S(p)\geq\tau_{R}^{s}\;\;\text{and}\;\;V(p)\geq\tau_{R}^{v},$ (14)
@@
Appendix D Discussion
D.1 Comparison with related work
Distinction from anchor-based approaches.
Although both CompMarkGS and SecureGS use an anchor-based structure, they are fundamentally different, starting from their problem definition. SecureGS (Zhang et al., 2025b) is a steganography method aiming to hide large amounts of data, with invisibility as its key metric. In contrast, CompMarkGS is a digital watermarking method for proving ownership, which prioritizes robustness. Notably, the ”robustness” evaluation in SecureGS focuses on preserving the rendering quality of the hidden scene, whereas CompMarkGS aims for the recovery of the hidden message itself under attacks like model compression. To achieve these different goals, their information embedding mechanisms are also distinct. SecureGS predicts hidden Gaussian offsets with a private MLP, while CompMarkGS directly injects a learnable watermark embedding feature into the anchor feature and extracts the message from the rendered image via a pre-trained decoder. Therefore, CompMarkGS is not an incremental improvement on SecureGS, but a unique solution independently designed for the new objective of verifiable ownership after model compression.
Comparison with 3DGS watermarking techniques.
A key difference between our method and existing watermarking methods, particularly frequency-based approaches such as 3D-GSW (Jang et al., 2025), lies in our unique approach to utilizing frequency information. While both methods leverage frequency data for Gaussian densification, their core strategies and objectives are fundamentally different. 3D-GSW measures the frequency intensity of patches, whereas CompMarkGS precisely identifies high-frequency error regions between the rendered image and the ground truth, selectively densifying anchors in those specific areas. However, the most significant distinction is the role of compression in the watermarking pipeline. Unlike 3D-GSW, which incorporates model pruning as part of its watermarking process, CompMarkGS is the first to define and solve the novel problem of watermark robustness against external quantization compression applied to a fully trained model. To achieve this, we introduce a quantization distortion layer (QDL) that simulates quantization noise during training, a core component absent in 3D-GSW. Our contribution is proposing a unique method specifically optimized to address the previously unexplored challenge of watermark loss during quantization compression. This focus on compression robustness also sets CompMarkGS apart from GaussianMarker (Huang et al., 2025) and GuardSplat (Chen et al., 2024b). These methods address traditional image distortions like Gaussian noise and JPEG compression, or model distortions such as pruning. However, they lack a mechanism to mitigate the loss of watermark information caused by quantization compression.
Furthermore, for a fair comparison, we adapt the existing vanilla 3DGS-based watermarking techniques to the anchor-based 3DGS framework. However, during this process, GuardSplat requires significant adaptation due to a core architectural incompatibility. This is because GuardSplat relies on adding offsets to explicit Spherical Harmonics (SH) parameters, which do not exist in our anchor-based model. Therefore, we modified its mechanism to apply offsets to the anchor features instead for our experiments.
D.2 Motivation for watermarking anchor-based 3D Gaussian Splatting
As the demand for lightweight and efficiently rendered 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) models grows, anchor-based architectures are emerging as a key solution. In particular, models like Scaffold-GS (Lu et al., 2024), which combine anchors with a MLP, achieve both high memory efficiency and excellent rendering quality compared to vanilla 3DGS. This anchor-based structure is highly practical, making it ideal for mobile and AR/VR environments where real-time rendering is essential.
We identified these anchor-based 3DGS models as ideal targets for watermarking due to a significant security advantage. Instead of directly storing and rendering Gaussian attributes, anchor-based models generate them indirectly by feeding anchor features into an MLP. If we embed a watermark into these anchor features, it undergoes a non-linear transformation by the MLP. This process expresses the watermark within the Gaussian attributes in a complex and concealed manner, making it extremely difficult for an external attacker to directly interpret or remove it.
Therefore, selecting an anchor-based 3DGS model for watermarking is a rational approach that secures both practical benefits, such as a lightweight structure, and robust security through the MLP.
Appendix E Additional experimental results
E.1 Effectiveness of watermarking embedding feature
To verify the effectiveness of our proposed watermark embedding strategy, we compare the performance of two approaches: using only the anchor feature and using a dedicated watermark embedding feature. As described in Sec.4.1, our method introduces a dedicated watermark embedding feature that shares the same dimension as the anchor feature, combining them via element-wise sum. This design ensures that the total feature dimension remains constant, regardless of whether a watermark is embedded. Therefore, any observed performance increase can be attributed to the effectiveness of the embedding strategy itself, rather than an increase in model capacity.
@@ E.1 Effectiveness of watermarking embedding feature
Table 7: Ablation study on the dedicated watermark embedding feature. Results are reported for 48-bit messages and averaged over three datasets after ContextGS compression.
@@
Embedding Method Bit Acc. (%) ↑ PSNR ↑ SSIM ↑ LPIPS ↓ Anchor Feature (Ours w/o $f^{\prime}$) 92.94 26.79 0.834 0.209 Anchor Feature (Ours w/ $f^{\prime}$) 94.03 27.55 0.844 0.173
As shown in Tab.7, using only the anchor feature creates a conflict between the two tasks of representing the scene and hiding information, which leads to degraded rendering quality. In contrast, our approach of introducing a dedicated watermark embedding feature preserves high rendering quality while enabling a more robust watermark. We conclude that using an independent watermark embedding feature is essential for effectively decoupling these two objectives, thereby achieving an optimal balance between rendering quality and watermark robustness.
E.2 Effectiveness of HSV loss